# PHANGS-JWST: Data Processing Pipeline and First Full Public Data Release

### v1

@@ 5.2 Comparison to MAST mosaics
Figure 15: Comparison between our reprocessing and Level Three MAST-obtained mosaics for NGC 628. The difference here is the MAST version subtracted from our reprocessed version, after reprojecting to the same WCS. As can be seen, the astrometry is somewhat different in NIRCam, and the striping in the MAST mosaic is clear. The background levels are substantially different in the MIRI bands, due to a combination of our improved background matching algorithm and the fact that the MAST mosaics do not include our dedicated background observations.
Here, we make a comparison to the Level Three mosaics that are available from MAST [^12] As downloaded on 2023/12/06. (https://mast.stsci.edu/). These images are automatically processed using observatory-recommended default settings per-band, and as such are designed to produce reasonable quality mosaics across the full breadth of science that JWST will cover.
@@ 6 Released Data Products
Extension name
Description
SCI
2-D data array containing the pixel values, in units of surface brightness (MJy/sr)
ERR
2-D data array containing uncertainty estimates for each pixel. These values are based on the combined VAR_POISSON and VAR_RNOISE data, given as standard deviation
DQ
2-D data array containing data quality flags for each pixel
VAR_POISSON
2-D data array containing the variance estimate for each pixel, based on Poisson noise only
VAR_RNOISE
2-D data array containing the variance estimate for each pixel, based on read noise only
VAR_FLAT
2-D data array containing the variance estimate for each pixel, based on uncertainty in the flat-field
[…]
3-D context image, which encodes information about which input images contribute to a specific output pixel
[…]
2-D weight image giving the relative weight of the output pixels
Table 2: Description of fits extensions for our released data products. ${}^{1}$nly present in img.fits files.
Our data products are hosted at the MAST high-level science products (HLSP) webpages, at https://archive.stsci.edu/hlsp/phangs-jwst, as well as on CANFAR at https://www.canfar.net/storage/vault/list/phangs/RELEASES/PHANGS-JWST. Our primary data products are mosaics (img.fits), although we also provide compressed directories containing individually corrected tiles (tweakback.tar.gz), which are fully processed up to the final mosaic stage, and have absolute astrometric corrections back-propagated to them. The files are presented as multi-extension .fits files, where the extensions are standard outputs from the JWST pipeline. These are detailed in Table 2. We note that pjpipe does not currently propagate the effects of any convolution to any arrays except the science and error arrays, and so these other extensions will be invalid for PSF matched images.
@@ 7 Summary
In this work, we have presented the overview of the steps we have taken in producing science-ready data for the PHANGS-JWST survey, implemented in our custom pipeline pjpipe (our DR1.0.1 was produced using pjpipe version 1.0.2). This pipeline provides:
@@
We have extensively tested our processing methodology for 19 galaxies observed in four NIRCam and four MIRI bands, and our approach produces excellent mosaics across our sample. Further testing on the 55 galaxies that make up our Cycle 2 Treasury (program 3707, PI: Leroy) will help refine the data processing further. The automated pipeline enables rapid processing of these data so that we can make products available in a timely manner. We have developed a number of bespoke algorithms that will be of general use to the community, and hence our pipeline, pjpipe, is made public at https://github.com/phangsTeam/pjpipe, and is pip-installable (pip install pjpipe). We have designed pjpipe to be flexible and easy to interface with, to maximize its usefulness. We make our first full public data release (DR1.0.1) mosaics available at https://archive.stsci.edu/hlsp/phangs-jwst, and stress that these represent a significant improvement on those we used in our early science papers, as well as those obtainable from the MAST database.
@@
This work has been carried out as part of the PHANGS collaboration. This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 2107. The specific observations analyzed can be accessed via doi:10.17909/ew88-jt15. TGW thanks P. Woolford for technical assistance throughout the preparation of this work. We also thank M. García Marín for comments on the manuscript. KS, JS acknowledge funding from JWST-GO-2107.006-A. R.S.K. and S.C.O.G. acknowledge funding from the European Research Council via the Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). AKL, DP, and RC gratefully acknowledges support by grants 1653300 and 2205628 from the National Science Foundation, by award JWST-GO-02107.009-A, JWST-GO-03707.001-A, JWST-GO-04256.001-A, and by a Humboldt Research Award from the Alexander von Humboldt Foundation. MC gratefully acknowledges funding from the DFG through an Emmy Noether Research Group (grant number CH2137/1-1). COOL Research DAO is a Decentralized Autonomous Organization supporting research in astrophysics aimed at uncovering our cosmic origins. J.K. is supported by a Kavli Fellowship at the Kavli Institute for Particle Astrophysics and Cosmology (KIPAC). JPe acknowledges support by the French Agence Nationale de la Recherche through the DAOISM grant ANR-21-CE31-0010 and by the Programme National “Physique et Chimie du Milieu Interstellaire” (PCMI) of CNRS/INSU with INC/INP, co-funded by CEA and CNES. OE acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the form of an Emmy Noether Research Group (grant number KR4598/2-1, PI Kreckel). JDH gratefully acknowledges financial support from the Royal Society (University Research Fellowship; URF/R1/221620). ES acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 694343)

### v3

@@ 5.2 Comparison to MAST mosaics
Figure 15: Comparison between our reprocessing and Level Three MAST-obtained mosaics for NGC 628. The difference here is the MAST version subtracted from our reprocessed version, after reprojecting to the same WCS. As can be seen, the astrometry is somewhat different in NIRCam, and the striping in the MAST mosaic is clear. The background levels are substantially different in the MIRI bands, due to a combination of our improved background matching algorithm and the fact that the MAST mosaics do not include our dedicated background observations. The complete figure set (19 images) is available in the online journal.
Here, we make a comparison to the Level Three mosaics that are available from MAST [^13] As downloaded on 2024/04/02. (https://mast.stsci.edu/). These images are automatically processed using observatory-recommended default settings per-band, and as such are designed to produce reasonable quality mosaics across the full breadth of science that JWST will cover.
@@ 6 Released Data Products
Extension name Description
SCI 2-D data array containing the pixel values, in units of surface brightness (MJy/sr)
ERR 2-D data array containing uncertainty estimates for each pixel. These values are based on the combined VAR_POISSON and VAR_RNOISE data, given as standard deviation
DQ 2-D data array containing data quality flags for each pixel
VAR_POISSON 2-D data array containing the variance estimate for each pixel, based on Poisson noise only
VAR_RNOISE 2-D data array containing the variance estimate for each pixel, based on read noise only
VAR_FLAT 2-D data array containing the variance estimate for each pixel, based on uncertainty in the flat-field
CON1 3-D context image, which encodes information about which input images contribute to a specific output pixel
WHT1 2-D weight image giving the relative weight of the output pixels
Table 3: Description of fits extensions for our released data products. 1Only present in img.fits files.
Our data products are hosted at the MAST high-level science products (HLSP) webpages, at https://archive.stsci.edu/hlsp/phangs/phangs-jwst, as well as on CANFAR at https://www.canfar.net/storage/vault/list/phangs/RELEASES/PHANGS-JWST. Our primary data products are mosaics (img.fits), although we also provide compressed directories containing individually corrected tiles (tweakback.tar.gz), which are fully processed up to the final mosaic stage, and have absolute astrometric corrections back-propagated to them. The files are presented as multi-extension .fits files, where the extensions are standard outputs from the JWST pipeline. These are detailed in Table 3. We note that pjpipe does not currently propagate the effects of any convolution to any arrays except the science and error arrays, and so these other extensions will be invalid for PSF matched images.
@@ 7 Summary
In this work, we have presented the overview of the steps we have taken in producing science-ready data for the PHANGS-JWST survey, implemented in our custom pipeline pjpipe (our DR1.1.0 was produced using pjpipe version 1.1.0). This pipeline provides:
@@
We have extensively tested our processing methodology for 19 galaxies observed in four NIRCam and four MIRI bands, and our approach produces excellent mosaics across our sample. Further testing on the 55 galaxies that make up our Cycle 2 Treasury (program 3707, PI: Leroy) will help refine the data processing further. The automated pipeline enables rapid processing of these data so that we can make products available in a timely manner. We have developed a number of bespoke algorithms that will be of general use to the community, and hence our pipeline, pjpipe, is made public at https://github.com/phangsTeam/pjpipe, and is pip-installable (pip install pjpipe). We have designed pjpipe to be flexible and easy to interface with, to maximize its usefulness. We make our first full public data release (DR1.1.0) mosaics available at https://archive.stsci.edu/hlsp/phangs/phangs-jwst, and stress that these represent a significant improvement on those we used in our early science papers, as well as those obtainable from the MAST database.
@@
This work has been carried out as part of the PHANGS collaboration. This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 2107. The specific observations analyzed can be accessed via doi:10.17909/ew88-jt15. The authors would like to thank the anonymous referee for their constructive comments, which have improved this work. We also thank M. García Marín for comments on the manuscript. TGW would like to personally thank P. Woolford for technical assistance throughout the preparation of this work, as well as Beryl Williams for everything throughout the years. KS, JS acknowledge funding from JWST-GO-2107.006-A. R.S.K. and S.C.O.G. acknowledge funding from the European Research Council via the Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). AKL, DP, and RC gratefully acknowledges support by grants 1653300 and 2205628 from the National Science Foundation, by award JWST-GO-02107.009-A, JWST-GO-03707.001-A, JWST-GO-04256.001-A, and by a Humboldt Research Award from the Alexander von Humboldt Foundation. MC gratefully acknowledges funding from the DFG through an Emmy Noether Research Group (grant number CH2137/1-1). COOL Research DAO is a Decentralized Autonomous Organization supporting research in astrophysics aimed at uncovering our cosmic origins. J.K. is supported by a Kavli Fellowship at the Kavli Institute for Particle Astrophysics and Cosmology (KIPAC). JPe acknowledges support by the French Agence Nationale de la Recherche through the DAOISM grant ANR-21-CE31-0010 and by the Programme National “Physique et Chimie du Milieu Interstellaire” (PCMI) of CNRS/INSU with INC/INP, co-funded by CEA and CNES. OE acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the form of an Emmy Noether Research Group (grant number KR4598/2-1, PI Kreckel). JDH gratefully acknowledges financial support from the Royal Society (University Research Fellowship; URF/R1/221620). ES acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 694343). SD is supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement no. 101018897 CosmicExplorer).